[Hallc_running] [New Logentry] RC daily update

simona at jlab.org simona at jlab.org
Tue Jul 16 11:25:03 EDT 2019

Logentry Text:
During day shift yesterday OPS searched for RF instabilities and found that 1L06-5 had a bad tuner stepper card. The card was replaced and the cavity was brought back online. After this Hall C received up to 70uA CW with Hall A running at low current for Moller commissioning. This morning at the end of the owl shift when the Hall A Moller commissioning ended, both Hall A and C received high current at the same time for a little while (https://logbooks.jlab.org/entry/3704494 , https://logbooks.jlab.org/entry/3704927 , https://logbooks.jlab.org/entry/3704937).

This morning at 10 am we finished data taking at 4 pass. Hall C will go to controlled access to set HMS to small angle (10.96 deg) in preparation for 5 pass running (a different beast).

There will be an RC meeting at 4:30 pm on the second floor of the counting house: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3704963
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190716/25aa9c46/attachment.html>

More information about the Hallc_running mailing list